Nice Corporation produces and sells a single product. Data concerning that product appear below:

Per Unit Percent of Sales Selling price $240 100 %

Variable expenses 48 20 %

Contribution margin $192 80%

Fixed expenses are $160,000 per month. The company is currently selling 1,500 units per month.

Required: Management is considering using a new component that would increase the unit variable cost by $65. Since the new component would improve the company's product, the marketing manager predicts that monthly sales would increase by 700 units. What should be the overall effect on the company's monthly net operating income of this change if fixed expenses are unaffected?
Answer #1

Current Net income for a month:

Contribution margin = 1500*192 = $288,000

Fixed cost = $160,000

Net Income = Contribution margin - Fixed cost

   = $288000-$160,000 = $128,000

Case II - Sales increased by 700 units and variable cost increased by $65 i.e(48+65 = 113 per unit variable cost)

Contribution margin = (1500+700)*(240-48-65) = 279400

Fixed cost = $160,000

Net Income = Contribution margin - Fixed cost

   = $279400-$160,000 = $119,400

Change : Monthly income will decreased by 128000-1194000 = 8600.